In a triangle ABC, prove that c.cos(A-x)+acos(C+x) = bcosx

Grade:11

1 Answers

Lab Bhattacharjee
8 years ago
HINT:   \text{Expand using }\cos(A\pm B) \text{ and use }b=a\cos C+c\cos A